Let’s say that you want to make $1 000 000 through regular investments. How much would you need to contribute each year if:
The interest rate was 2% and you invested for twenty years?
The interest rate was 4% and you invested for ten years?
The interest rate was 4% and you invested for twenty years?
The interest rate was 10% and you invested for twenty years?
1
Expert's answer
2019-09-24T11:22:40-0400

The formula for ordinary annuity is:

P = r×FV/((1 + r)^n - 1)

1) P = 0.02×1,000,000/((1 + 0.02)^(20) - 1) = 41,156.71,

2) P = 0.04×1,000,000/((1 + 0.04)^(10) - 1) = 83,290.94,

3) P = 0.04×1,000,000/((1 + 0.04)^(20) - 1) = 33,581.75,

4) P = 0.1×1,000,000/((1 + 0.1)^(20) - 1) = 17,459.62.